Perform the indicated operation. Let f(x)=x+1 and g(x)=x-1. Find f(g(x)).: Perform the indicated operation. Let f(x)=x+1 and g(x)=x-1. Find f(g(x)).
Answer by jim_thompson5910(9869)

f(x)=x+1 Start with the first function


f(g(x))=(x-1)+1 Plug in g(x)=x-1


f(g(x))=x-1+1 Remove the parenthesis


f(g(x))=x Combine like terms.